# An Introduction to Reinforcement Learning: Fundamental Concepts and Practical Applications

### v1

@@ 3 Finite Markov Decision Process
This section describes the formal problem of finite Markov decision processes (MDPs). An MDP provides a framework for sequential decision-making in which actions affect immediate rewards as well as future outcomes. In MDPs, immediate rewards are balanced with delayed rewards. In contrast to bandit problems in which the goal is to determine the value of each action $a$, MDPs aim to measure the value of taking action $a$ in state $s$, or the value of being in state $s$ assuming optimal actions are taken. A correct assessment of the long-term effects of interventions requires the estimation of these state-specific values [sutton2018reinforcement]. Finite MDPs consist of states, actions, and rewards $(S,A,R)$. Discrete probability distributions are assigned to the random variables $R_{t}$ and $S_{t}$ based on the preceding state and action. Using the probabilities of occurrence of the random variables $R_{t}$ and $S_{t}$, derive equations for these variables. A system is considered Markovian when the outcome of an action is independent of past actions and states, relying solely on the current state [van2012reinforcement]. The Markov property requires the state to encapsulate significant details of the entire past interaction influencing future outcomes [kaelbling1996reinforcement]. This definition is the basis of MDPs being used in RL. To describe the dynamics of an MDP, we use the state-transition probability function $p(s^{\prime},r\mid s,a)$, which is defined as follows:
$p(s^{\prime},r\mid s,a)\equiv\Pr\{S_{t}=s^{\prime},R_{t}=r\mid S_{t-1}=s,A_{t-1}=a\}$ (12)
where the function $p$ defines the MDP dynamics.
@@
The following state-transition probabilities, state-action and state-action-next-state triple rewards can be derived from the four-argument dynamic function $p$. We can derive the state-transition probabilities, the expected reward for state-action pairs, and the expected rewards for state-action-next-state triples as follows:
$p(s^{\prime}\mid s,a)\equiv\Pr\{S_{t}=s^{\prime}\mid S_{t-1}=s,A_{t-1}=a\}=\sum_{r\in R}p(s^{\prime},r\mid s,a)$ (13)
$r(s,a)\equiv\mathbb{E}\{R_{t}\mid S_{t-1}=s,A_{t-1}=a\}=\sum_{r\in R}r\sum_{r\in R}p(s^{\prime},r\mid s,a)$ (14)
$r(s,a,s^{\prime})\equiv\mathbb{E}\{R_{t}\mid S_{t-1}=s,A_{t-1}=a,S_{t}=s^{\prime}\}=\sum_{r\in R}\frac{rp(s^{\prime},r\mid s,a)}{p(s^{\prime}\mid s,a)}$ (15)
@@
The concept of actions encompasses any decisions relating to learning, and the concept of states encompasses any information that is available in order to inform those decisions. As part of the MDP framework, goal-directed behavior is abstracted through interaction. Any learning problem can be reduced to three signals between an agent and its environment: actions, states, and rewards. A wide range of applications have been demonstrated for this framework [sutton2018reinforcement].
We are now able to formally define and solve RL problems. We have defined rewards, objectives, probability distributions, the environment, and the agent. Some concepts, however, were defined informally. According to our statement, the agent seeks to maximize future rewards, but how can this be mathematically expressed?
The return, denoted $G_{t}$, is the cumulative sum of rewards received from time step $t$ onwards. For episodic tasks, it is defined as follows:
Let the sequence of rewards received after time step $t$ be:
$G_{t}\equiv R_{t+1}+R_{t+2}+\ldots+R_{T}$ (16)
Here, $G_{t}$ is a specific function of the reward sequence.
In what way does terminating the sequence at $T$ serve a purpose? As the name suggests, episodic problems are those in which the interactions between agents and their environment occur naturally in sequence, known as episodes, and tasks are termed episodic tasks. The game hangman is a good example of this. At the end of each episode, a standard starting state is restored. The term ”new games” refers to the next state after the terminal state, which is the final state leading to the end of an episode.
@@
It is common for ongoing tasks to involve interactions that persist continuously throughout the duration of the task, such as process control or applications that utilize robots with prolonged lifespans. The term ”continuing tasks” refers to these activities. As there are no terminal states in continuing tasks $(T=/infty)$, the return for continuing tasks should be defined differently. It is possible that the return could be infinite if the agent consistently receives a reward. For continuing tasks, define the return as follows:
For continuing tasks, where there is no terminal state, the return $G_{t}$ is defined as the discounted sum of future rewards:
$G_{t}\equiv R_{t+1}+\gamma R_{t+2}+\gamma^{2}R_{t+3}+\ldots=\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}$ (17)
where $\gamma$ is the discount rate $(0\leq\gamma\leq 1)$.
@@
The discount rate affects the current worth of future rewards. When $\gamma<1$, the infinite sum converges to a finite value. With $\gamma=0$, the agent maximizes immediate rewards. As $\gamma$ approaches 1, future rewards carry more weight. Define return recursively as:
We can also express the return $G_{t}$ recursively as:
$G_{t}\equiv R_{t+1}+\gamma G_{t+1}$ (18)
@@
The return is finite if the reward is nonzero and constant, and $\gamma<1$. Use one formula for both episodic and continuing tasks:
For both episodic and continuing tasks, the return $G_{t}$ can be defined as:
$G_{t}\equiv\sum_{k=t+1}^{T}\gamma^{k-t-1}R_{k}$ (19)
This formula works for both episodic and continuing tasks if $T=\infty$ or $\gamma=1$, respectively.
Now, we can discuss the two most important RL concepts: Policies and Value functions.
4 Policies and Value Functions: Zero to Hero
@@
The value function estimates the expected return of the agent being in a certain state (or performing an action in a particular state). Depending on the actions selected, these factors will vary. There is a link between value functions and policies, which are linked to probabilities of action based on states. Value functions can be divided into two broad categories as follows:
State Value Functions: The value function of a state $s$ under policy $\pi$, $v_{\pi}(s)$, is the expected return starting in $s$ and following $\pi$ thereafter.
Action Value Functions: The value of taking action $a$ in state $s$ under policy $\pi$, $q_{\pi}(s,a)$, is the expected return starting from $s$, taking action $a$, and following $\pi$ thereafter.
For MDPs, $v$ and $q$ are defined as:
The state value function $v_{\pi}(s)$ represents the expected return starting from state $s$ and following policy $\pi$. It is mathematically defined as follows:
@@
$v_{\pi}(s)\equiv\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\mid S_{t}=s\right],\quad\text{for all }s\in S$ (20)
The action value function $q_{\pi}(s,a)$ represents the expected return starting from state $s$, taking action $a$, and then following policy $\pi$. It is defined as follows:
$q_{\pi}(s,a)\equiv\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s,A_{t}=a]=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\mid S_{t}=s,A_{t}=a\right]$ (21)
@@
It is important to note the difference between $v$ and $q$, namely that $q$ depends on the actions taken in each state. With ten states and eight actions per state, $q$ requires 80 functions, while $v$ requires only 10 functions.
Following policy $\pi$, if an agent averages returns from each state, the average converges to $v^{\pi}(s)$. Averaging returns from each action converges to $q^{\pi}(s,a)$ [sutton2018reinforcement].
In Monte Carlo methods, many random samples of returns are averaged. This method does not provide sample efficiency, requiring separate averages to be calculated for each state. Estimates can be improved through the use of parameterized functions that have fewer parameters. $v$ should be written recursively as follows:
$v_{\pi}(s)\equiv\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s]=\mathbb{E}_{\pi}[R_{t+1}+\gamma G_{t+1}\mid S_{t}=s]=\sum_{a}\pi(a|s)\sum_{s^{\prime}}\sum_{r}p(s^{\prime},r|s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (22)
@@
This is the Bellman equation for $v_{\pi}$. The Bellman equation relates the value of a state to its potential successor states’ values. The diagram illustrates the anticipation from a state to its successors. The value of the initial state equals the discounted value of the expected next state plus the anticipated reward [o2018uncertainty, sutton2018reinforcement].
4.1 When to Use What?
State-value functions $v_{\pi}(s)$ and action-value functions $q_{\pi}(s,a)$ serve different purposes in RL.
In the evaluation of deterministic policies or when understanding the value of being in a particular state is required, state-value functions are used. In policy evaluation and policy iteration methods, where a policy is explicitly defined and it is necessary to evaluate the performance of being in a particular state under the policy, these methods are highly useful. The use of state-value functions is beneficial when there are many actions, since they reduce complexity by requiring only an evaluation of state values.
@@ 4.1 When to Use What?
Action-value functions are used to evaluate and compare the potential for different actions when they are taking place in the same state. They are crucial for the selection of actions, such as in Q-learning and SARSA, where the goal is to determine the most appropriate action for each situation. As action-value functions take into account the expected return of different actions, they are particularly useful in environments with stochastic policies. Moreover, when dealing with continuous action spaces, action-value functions can provide a more detailed understanding of the impact of actions, aiding in the fine-tuning of policy implementation.
@@
Consider a gambling scenario where a player starts with \$10 and faces decisions regarding the amount to bet. This game illustrates state and action value functions in RL.
State Value function ($v^{\pi}(s)$): The state value function $v^{\pi}(s)$ quantifies expected cumulative future rewards for a state $s$, given policy $\pi$. Suppose the player has \$5:
  • With a consistent \$1 bet, $v^{\pi}(5)=0.5$ indicates an expected gain of \$0.5.
  • With a consistent \$2 bet, $v^{\pi}(5)=-1$ indicates an expected loss of \$1.
@@
Action Value function ($q^{\pi}(s,a)$): The action value function $q^{\pi}(s,a)$ assesses expected cumulative future rewards for action $a$ in state $s$. For instance:
  • $q^{\pi}(5,1)=1$ suggests a \$1 bet from \$5 results in a cumulative reward of \$1.
  • $q^{\pi}(5,2)=-0.5$ indicates a loss of \$0.5 for a \$2 bet from \$5.
@@
Solving an RL task involves identifying a policy that maximizes long-term rewards. For finite MDPs, an optimal policy is defined in the next subsection.
@@ 4.2 Optimal Policies and Optimal Value Functions
Value functions create a partial ordering over policies, allowing comparison and ranking based on expected cumulative rewards. A policy $\pi$ is better than or equal to $\pi_{0}$ if $v^{\pi}(s)\geq v^{\pi_{0}}(s)$ for all states $s$. An optimal policy is better than or equal to all other policies, denoted by $\pi_{*}$, sharing the same optimal state-value function $v^{*}$:
The optimal state value function $v^{*}(s)$ is defined as the maximum value function over all possible policies:
$v^{*}(s)\equiv\max_{\pi}v^{\pi}(s)\quad\text{for all }s\in S$ (23)
Optimal policies also share the same optimal action-value function $q^{*}$:
The optimal action value function $q^{*}(s,a)$ is defined as the maximum action value function over all possible policies:
$q^{*}(s,a)\equiv\max_{\pi}q^{\pi}(s,a)\quad\text{for all }s\in S$ (24)
The relationship between the optimal action value function $q^{*}(s,a)$ and the optimal state value function $v^{*}(s)$ is given by the following equation:
$q^{*}(s,a)=\mathbb{E}[R_{t+1}+\gamma v^{*}(S_{t+1})\mid S_{t}=s,A_{t}=a]$ (25)
The equation expresses expected cumulative return for state-action pairs in terms of immediate rewards and discounted future states’ values.
Optimal value functions and policies represent an ideal state in RL. It is however rare to find truly optimal policies in computationally demanding tasks due to practical challenges. RL agents strive to approximate optimal policies. Dynamic Programming (DP) helps identify optimal values, assuming a perfect model of the environment.
The fundamental idea of DP and RL is using value functions to organize the search for good policies. For finite MDPs, the environment’s dynamics are given by probabilities $p(s^{\prime},r\mid s,a)$. DP finds exact solutions in special cases, like finding the shortest path in a graph.
The Bellman optimality equations for the optimal state value function $v^{*}(s)$ and the optimal action value function $q^{*}(s,a)$ are given as follows:
$v^{*}(s)=\max_{a}\mathbb{E}[R_{t+1}+\gamma v^{*}(S_{t+1})\mid S_{t}=s,A_{t}=a]=\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v^{*}(s^{\prime})]$ (26)
$q^{*}(s,a)=\mathbb{E}[R_{t+1}+\max_{a^{\prime}}q^{*}(S_{t+1},a^{\prime})\mid S_{t}=s,A_{t}=a]=\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma\max_{a^{\prime}}q^{*}(s^{\prime},a^{\prime})]$ (27)

### v2

@@ 2.2 Markov Decision Process (MDPs)
An MDP provides a framework for sequential decision-making in which actions affect immediate rewards as well as future outcomes. In MDPs, immediate rewards are balanced with delayed rewards. In contrast to bandit problems in which the goal is to determine the value of each action $a$, MDPs aim to measure the value of taking action $a$ in state $s$, or the value of being in state $s$ assuming optimal actions are taken. A correct assessment of the long-term effects of interventions requires the estimation of these state-specific values [1]. MDPs consist of states, actions, and rewards $(\mathcal{S,A,R})$. Discrete probability distributions are assigned to the random variables $\mathcal{R}_{t}$ and $\mathcal{S}_{t}$ based on the preceding state and action. Using the probabilities of occurrence of the random variables $\mathcal{R}_{t}$ and $\mathcal{S}_{t}$, derive equations for these variables. A system is considered Markovian when the outcome of an action is independent of past actions and states, relying solely on the current state [9]. The Markov property requires the state to encapsulate significant details of the entire past interaction influencing future outcomes [10]. This definition is the basis of MDPs being used in RL. To describe the dynamics of an MDP, we use the state-transition probability function $p(s^{\prime},r\mid s,a)$, which is defined as follows:
$p(s^{\prime},r\mid s,a)\equiv\Pr\{\mathcal{S}_{t}=s^{\prime},\mathcal{R}_{t}=r\mid\mathcal{S}_{t-1}=s,\mathcal{A}_{t-1}=a\}$ (9)
@@
where the function $p$ defines the MDP dynamics. The following state-transition probabilities, state-action and state-action-next-state triple rewards can be derived from the four-argument dynamic function $p$. We can derive the state-transition probabilities, the expected reward for state-action pairs, and the expected rewards for state-action-next-state triples as follows:
$p(s^{\prime}\mid s,a)\equiv\Pr\{\mathcal{S}_{t}=s^{\prime}\mid\mathcal{S}_{t-1}=s,\mathcal{A}_{t-1}=a\}=\sum_{r\in\mathcal{R}}p(s^{\prime},r\mid s,a)$ (10)
$r(s,a)\equiv\mathbb{E}\{\mathcal{R}_{t}\mid\mathcal{S}_{t-1}=s,\mathcal{A}_{t-1}=a\}=\sum_{r\in\mathcal{R}}r\sum_{r\in\mathcal{R}}p(s^{\prime},r\mid s,a)$ (11)
$r(s,a,s^{\prime})\equiv\mathbb{E}\{\mathcal{R}_{t}\mid\mathcal{S}_{t-1}=s,\mathcal{A}_{t-1}=a,\mathcal{S}_{t}=s^{\prime}\}=\sum_{r\in\mathcal{R}}\frac{rp(s^{\prime},r\mid s,a)}{p(s^{\prime}\mid s,a)}$ (12)
@@
The concept of actions encompasses any decisions relating to learning, and the concept of states encompasses any information that is available in order to inform those decisions. As part of the MDP framework, goal-directed behavior is abstracted through interaction. Any learning problem can be reduced to three signals between an agent and its environment: actions, states, and rewards. A wide range of applications have been demonstrated for this framework [1]. We are now able to formally define and solve RL problems. We have defined rewards, objectives, probability distributions, the environment, and the agent. Some concepts, however, were defined informally. According to our statement, the agent seeks to maximize future rewards, but how can this be mathematically expressed? The return, denoted $G_{t}$, is the cumulative sum of rewards received from time step $t$ onwards. For episodic tasks, it is defined as follows:
$\mathcal{G}_{t}\equiv\mathcal{R}_{t+1}+\mathcal{R}_{t+2}+\ldots+\mathcal{R}_{\mathcal{T}}$ (13)
@@
Here, $\mathcal{G}_{t}$ is a specific function of the reward sequence. Episodic problems are those in which the interactions between agents and their environment occur naturally in sequence, known as episodes, and tasks are termed episodic tasks. The game hangman is a good example of this. At the end of each episode, a standard starting state is restored. The term new games refers to the next state after the terminal state, which is the final state leading to the end of an episode. It is common for ongoing tasks to involve interactions that persist continuously throughout the duration of the task, such as process control or applications that utilize robots with prolonged lifespans. The term continuing tasks refers to these activities. As there are no terminal states in continuing tasks $(\mathcal{T}=\infty)$, the return for continuing tasks should be defined differently. It is possible that the return could be infinite if the agent consistently receives a reward. For continuing tasks, where there is no terminal state, the return $\mathcal{G}_{t}$ is defined as the discounted sum of future rewards:
$\mathcal{G}_{t}\equiv\mathcal{R}_{t+1}+\gamma\mathcal{R}_{t+2}+\gamma^{2}\mathcal{R}_{t+3}+\ldots=\sum_{k=0}^{\infty}\gamma^{k}\mathcal{R}_{t+k+1}$ (14)
@@
where $\gamma$ is the discount rate $(0\leq\gamma\leq 1)$. The discount rate affects the current worth of future rewards. When $\gamma<1$, the infinite sum converges to a finite value. With $\gamma=0$, the agent maximizes immediate rewards. As $\gamma$ approaches 1, future rewards carry more weight. We can also express the return $\mathcal{G}_{t}$ recursively
$\mathcal{G}_{t}\equiv\mathcal{R}_{t+1}+\gamma\mathcal{G}_{t+1}$ (15)
@@
The return is finite if the reward is non-zero and constant, and $\gamma<1$. Equation 16 works for both episodic and continuing tasks if $\mathcal{T}=\infty$ or $\gamma=1$, respectively.
$\mathcal{G}_{t}\equiv\sum_{k=t+1}^{\mathcal{T}}\gamma^{k-t-1}\mathcal{R}_{k}$ (16)
The concepts introduced in this section—ranging from bandits and MDPs to value functions and policies—provide the mathematical and conceptual tools required to understand RL methods. In the next subsection, we explore how these foundational ideas evolve into core RL algorithms, bridging theory and application.
2.3 Policies and Value Functions
@@
The value function estimates the expected return of the agent being in a certain state (or performing an action in a particular state). Depending on the actions selected, these factors will vary. There is a link between value functions and policies, which are linked to probabilities of action based on states. Value functions can be divided into two broad categories; State Value Functions and Action Value Functions. The value function of a state $s$ under policy $\pi$, $v_{\pi}(s)$, is the expected return starting in $s$ and following $\pi$ thereafter (Equation 17). On the other hand, the value of taking action $a$ in state $s$ under policy $\pi$, $q_{\pi}(s,a)$, is the expected return starting from $s$, taking action $a$, and following $\pi$ thereafter (Equation 18).
@@
$v_{\pi}(s)\equiv\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}\mathcal{R}_{t+k+1}\mid\mathcal{S}_{t}=s\right],\quad\text{for all }s\in\mathcal{S}$ (17)
$q_{\pi}(s,a)\equiv\mathbb{E}_{\pi}[\mathcal{G}_{t}\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}\mathcal{R}_{t+k+1}\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a\right]$ (18)
@@
It is important to note the difference between $v$ and $q$, namely that $q$ depends on the actions taken in each state. With ten states and eight actions per state, $q$ requires 80 functions, while $v$ requires only 10 functions. Following policy $\pi$, if an agent averages returns from each state, the average converges to $v_{\pi}(s)$. Averaging returns from each action converges to $q_{\pi}(s,a)$ [1]. $v_{\pi}(s)$ can be written recursively:
$v_{\pi}(s)\equiv\mathbb{E}_{\pi}[\mathcal{G}_{t}\mid\mathcal{S}_{t}=s]=\mathbb{E}_{\pi}[\mathcal{R}_{t+1}+\gamma\mathcal{G}_{t+1}\mid\mathcal{S}_{t}=s]=\sum_{a}\pi(a|s)\sum_{s^{\prime}}\sum_{r}p(s^{\prime},r|s,a)[r+\gamma v_{\pi}(s^{\prime})]$ (19)
@@
Equation 19 is the Bellman equation for $v_{\pi}$. The Bellman equation relates the value of a state to its potential successor states’ values. The diagram illustrates the anticipation from a state to its successors. The value of the initial state equals the discounted value of the expected next state plus the anticipated reward [11, 1].
@@
$v_{\pi}(s)$ and $q_{\pi}(s,a)$ serve different purposes in RL. In the evaluation of deterministic policies or when understanding the value of being in a particular state is required, state-value functions are used. In policy evaluation and policy iteration methods, where a policy is explicitly defined and it is necessary to evaluate the performance of being in a particular state under the policy, these methods are highly useful. The use of state-value functions is beneficial when there are many actions, since they reduce complexity by requiring only an evaluation of state values. Action-value functions, on the other hand, are used to evaluate and compare the potential for different actions when they are taking place in the same state. They are crucial for the selection of actions, where the goal is to determine the most appropriate action for each situation. As action-value functions take into account the expected return of different actions, they are particularly useful in environments with stochastic policies. Moreover, when dealing with continuous action spaces, action-value functions can provide a more detailed understanding of the impact of actions, aiding in the fine-tuning of policy implementation.
@@
Example: Consider a gambling scenario where a player starts with \$10 and faces decisions regarding the amount to bet. This game illustrates state and action value functions in RL. State Value function ($v^{\pi}(s)$) quantifies expected cumulative future rewards for a state $s$, given policy $\pi$. Suppose the player has \$5:
  • With a consistent \$1 bet, $v_{\pi}(5)=0.5$ indicates an expected gain of \$0.5.
  • With a consistent \$2 bet, $v_{\pi}(5)=-1$ indicates an expected loss of \$1.
@@
Action Value function ($q_{\pi}(s,a)$) assesses expected cumulative future rewards for action $a$ in state $s$. For instance:
  • $q_{\pi}(5,1)=1$ suggests a \$1 bet from \$5 results in a cumulative reward of \$1.
  • $q_{\pi}(5,2)=-0.5$ indicates a loss of \$0.5 for a \$2 bet from \$5.
@@ 2.4 Optimal Policies and Optimal Value Functions
Solving an RL task involves identifying a policy that maximizes long-term rewards. Value functions create a partial ordering over policies, allowing comparison and ranking based on expected cumulative rewards. A policy $\pi$ is better than or equal to $\pi_{0}$ if $v_{\pi}(s)\geq v_{\pi_{0}}(s)$ for all states $s$. An optimal policy is better than or equal to all other policies, denoted by $\pi_{*}$, sharing the same optimal state-value function $v_{*}$, which is defined as the maximum value function over all possible policies.
$v_{*}(s)\equiv\max_{\pi}v_{\pi}(s)\quad\text{for all }s\in\mathcal{S}$ (20)
Optimal policies also share the same optimal action-value function $q_{*}$, which is defined as the maximum action value function over all possible policies.
$q_{*}(s,a)\equiv\max_{\pi}q_{\pi}(s,a)\quad\text{for all }s\in\mathcal{S}$ (21)
The relationship between the optimal action value function $q_{*}(s,a)$ and the optimal state value function $v_{*}(s)$ is given by the following equation: By having optimal action value function $q_{*}(s,a)$ we can find optimal state value function $v_{*}(s)$ as shown in Equation 22.
$q_{*}(s,a)=\mathbb{E}[\mathcal{R}_{t+1}+\gamma v_{*}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]$ (22)
Optimal value functions and policies represent an ideal state in RL. It is however rare to find truly optimal policies in computationally demanding tasks due to practical challenges [1]. RL agents strive to approximate optimal policies. Dynamic Programming (DP) helps identify optimal values, assuming a perfect model of the environment, which is a challenge to have in real-world cases. DP methods are not also sample efficient, even though they are theoretically sound. The fundamental idea of DP and RL is using value functions to organize the search for good policies. For finite MDPs, the environment’s dynamics are given by probabilities $p(s^{\prime},r\mid s,a)$. The Bellman optimality equations for the optimal state value function $v_{*}(s)$ and the optimal action value function $q_{*}(s,a)$ are Equations 23 and  24, respectively:
$v_{*}(s)=\max_{a}\mathbb{E}[\mathcal{R}_{t+1}+\gamma v_{*}(\mathcal{S}_{t+1})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\max_{a}\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma v_{*}(s^{\prime})]$ (23)
$q_{*}(s,a)=\mathbb{E}[\mathcal{R}_{t+1}+\max_{a^{\prime}}q_{*}(\mathcal{S}_{t+1},a^{\prime})\mid\mathcal{S}_{t}=s,\mathcal{A}_{t}=a]=\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r+\gamma\max_{a^{\prime}}q_{*}(s^{\prime},a^{\prime})]$ (24)